OpenStudy (anonymous):

Find the next three terms in the geometric sequence. Select 3 answers. 512, 384, 288, ...

OpenStudy (anonymous):

There is a pattern emerging when one sees that 512 = (3^0) x (2^9). 384 = (3^1) x (2^7). 288 = (3^2) x (2^5). Each term is [3^(n-1)] x [2^(11-2n)]. Do you see the same pattern?